# Fremlin tensor product respects the unbounded order convergence

### v2

@@ 3.1. Fremlin tensor product and $uo$-convergence
First, assume that $(f_{\alpha}),(g_{\alpha})\geq 0$. By considering [12, Lemma 3.1], we conclude that $f,g\geq 0$. Note that by the known Maeda-Ogasawara theorem ([1, Theorem 7.29]) and using Lemma 2, There are two compact Hausdorff extremally disconnected topological spaces $\Sigma$ and $\Omega$ such that $E$ is an order dense vector sublattice of $S(\Sigma)$ and $F$ is an order dense vector sublattice of $S(\Omega)$. Thus, by [1, Theorem 1.23], $E$ and $F$ are regular sublattices in $S(\Sigma)$ and $S(\Omega)$, respectively. So, by [11, Theorem 3.2], $f_{\alpha}\xrightarrow{uo}f$ in $S(\Sigma)$ and $g_{\alpha}\xrightarrow{uo}g$ in $S(\Omega)$. On the other hand, by [6, Proposition 3.1], $E\overline{\otimes}F$ can be considered as a vector sublattice of $S(\Sigma\times\Omega)$ generated by the mappings $(h\otimes k)(t,s)=h(t)k(s)$ for each $h\in S(\Sigma)$ and for each $k\in S(\Omega)$. Compatible with [5, Remark 4.1] and Lemma 2, we show that $f_{\alpha}\otimes g_{\alpha}\xrightarrow{uo}f\otimes g$ in $S(\Sigma\times\Omega)$. Note that the constant one function $\textbf{1}_{S(\Sigma)}$ is a weak unit for $S(\Sigma)$ and $\textbf{1}_{S(\Omega)}$ is a weak unit for $S(\Omega)$ so that $\textbf{1}_{S(\Sigma)}\otimes\textbf{1}_{S(\Omega)}$ is a weak unit for $S(\Sigma)\overline{\otimes}S(\Omega)$ by [3, Proposition 4.4] so that in $S(\Sigma\times\Omega)$ by Lemma 6. So, it is enough to consider $uo$-convergence just in terms of weak units by [11, Corollary 3.5]. Now, we use Lemma 5. Suppose $W\subseteq\Sigma\times\Omega$ is a non-empty open set and $\varepsilon\in(0,1)$ is arbitrary. There are some non-empty open sets $U\subseteq\Sigma$ and $V\subseteq\Omega$ such that $U\times V\subseteq W$. We can find non-empty open sets $U_{1}\subseteq U$ and $V_{1}\subseteq V$ with $(f_{\alpha}-f)(t)<\varepsilon$ for each $t\in U_{1}\cap Dom(f_{\alpha})\cap Dom(f)$ and $(g_{\alpha}-g)(s)<\varepsilon$ for each $s\in V_{1}\cap Dom(g_{\alpha})\cap Dom(g)$, provided that $\alpha$ is sufficiently large. Pick any $t_{0}$ in non-empty open set $U_{1}\cap Dom(f)$. Since $f$ is continuous at $t_{0}$ (WOLG, $f(t_{0})=1)$), there is a non-empty open set $U_{2}\subseteq U_{1}$ (containing $t_{0}$) such that $f(t)\leq 1$ for each $t\in U_{2}\cap Dom(f)$. Similarly, there exists a non-empty open set $V_{2}\subseteq V_{1}$ such that $g(s)\leq 1$ for each $s\in V_{2}\cap Dom(g)$. Then for each $(t,s)\in(U_{2}\times V_{2})\cap((Dom(f_{\alpha})\cap Dom(f))\times(Dom(g_{\alpha})\cap Dom(g)))=(U_{2}\cap Dom(f_{\alpha})\cap Dom(f))\times(V_{2}\cap Dom(g_{\alpha})\cap Dom(g)))$ and for sufficiently large $\alpha$, we have
$(f_{\alpha}\otimes g_{\alpha}-f\otimes g)(t,s)=(f_{\alpha}\otimes(g_{\alpha}-g))(t,s)+((f_{\alpha}-f)\otimes g)(t,s)$
$=f_{\alpha}(t)(g_{\alpha}-g)(s)+(f_{\alpha}-f)(t)g(s)<(\varepsilon+2)\varepsilon.$
@@
Now, an easy application of Lemma 6 and Lemma 7 convinces us that $f_{\alpha}\otimes g_{\alpha}\xrightarrow{uo}f\otimes g$ in $E\overline{\otimes}F$.
For the general case, observe that we can write $f_{\alpha}={f_{\alpha}}^{+}-{f_{\alpha}}^{-}$ and $g_{\alpha}={g_{\alpha}}^{+}-{g_{\alpha}}^{-}$. Moreover, by using [12, Lemma 3.1], we see that $f_{\alpha}\xrightarrow{uo}f$ if and only if ${f_{\alpha}}^{\pm}\xrightarrow{uo}{f}^{\pm}$ and also $g_{\alpha}\xrightarrow{uo}g$ if and only if ${g_{\alpha}}^{\pm}\xrightarrow{uo}{g}^{\pm}$. By using the former case, we have the following observation.
$f_{\alpha}\otimes g_{\alpha}=({f_{\alpha}}^{+}-{f_{\alpha}}^{-})\otimes({g_{\alpha}}^{+}-{g_{\alpha}}^{-})=({f_{\alpha}}^{+}\otimes{g_{\alpha}}^{+})-({f_{\alpha}}^{+}\otimes{g_{\alpha}}^{-})-({f_{\alpha}}^{-}\otimes{g_{\alpha}}^{+})+$
$({f_{\alpha}}^{-}\otimes{g_{\alpha}}^{-})\xrightarrow{uo}(f^{+}\otimes g^{+})-(f^{+}\otimes g^{-})-(f^{-}\otimes g^{+})+(f^{-}\otimes g^{-})=f\otimes g.$
@@
Suppose $E$ and $F$ are vector lattices. Assume that $(x_{\alpha})\subseteq E$ is $uo$-null and $(y_{\alpha})\subseteq F$ is eventually order bounded. Then, $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uo}0$ in $E\overline{\otimes}F$.
@@
There exists $y\in F_{+}$ such that $|y_{\alpha}|\leq y$ for sufficiently large $\alpha$. Suppose $w\in(E\overline{\otimes}F)_{+}$. By [10, 1A (d)]), there exist $x_{0}\in E_{+}$ and $y_{0}\in F_{+}$ with $w\leq x_{0}\otimes y_{0}$. Now, we use the following inequality.
$|x_{\alpha}\otimes y_{\alpha}|\wedge w\leq(|x_{\alpha}|\otimes|y_{\alpha}|)\wedge(x_{0}\otimes y_{0})$
$\leq(|x_{\alpha}|\wedge x_{0})\otimes(|y_{\alpha}|\vee y_{0})\leq(|x_{\alpha}|\wedge x_{0})\otimes(y\vee y_{0}).$
Now, by [13, Corollary 3.4], we conclude that $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uo}0$, as claimed. ∎
@@
Note that order bounded assumption is essential in Proposition 9 and can not be removed. It is know that in $L_{p}(\mu)$-spaces, for sequences, $uo$-convergence and almost every where convergence agree. Put $E=L^{2}(\mathbb{R})$, $f_{n}(x)=\frac{x}{n}$ and $g_{n}(x)=n$. It is easy to see that $(f_{n})$ is $uo$-null and $(g_{n})$ is not order bounded. Moreover, $(f_{n}\otimes g_{n})_{n}$ is the identity mapping which is not pointwise convergent to zero, certainly. More precisely, by [10, Proposition 4B], $L^{2}(X)\widehat{\otimes}L^{2}(Y)$ can be considered as vector sublattice of $L^{2}(X\times Y)$, where $X$ and $Y$ are measure spaces. Moreover, $L^{2}(X)\overline{\otimes}L^{2}(Y)$ is an order dense in $L^{2}(X)\widehat{\otimes}L^{2}(Y)$.
It can be easily verified that if a net $(x_{\alpha})\subseteq E$ and a net $(y_{\alpha})\subseteq F$ are eventually order bounded, then, the net $(x_{\alpha}\otimes y_{\alpha})$ is also eventually order bounded in $E\overline{\otimes}F$. So, we have the following fact that is an extension of [13, Corollary 3.4], as well.
@@
3.2. Fremlin projective tensor product and unbounded convergences
[…]
Lemma 12.
Suppose $E$ and $F$ are Banach lattices such that the Fremlin projective tensor product $E\widehat{\otimes}F$ is order continuous. Then, both $E$ and $F$ are order continuous.
Proof.
We show that $E$ is order continuous; order continuity of $F$ is similar. Suppose on a contrary, $E$ is not order continuous. So, there exists a disjoint order bounded sequence $(x_{n})\subseteq E$ which is not norm null. Choose any non-zero positive $y\in F$. It is easy to see that the sequence $(x_{n}\otimes y)$ is disjoint and order bounded in $E\widehat{\otimes}F$. By the assumption, it is norm null. But, the projective tensor norm is a cross-norm so that $(x_{n})$ should be norm null that is a contradiction. ∎
Observe that the converse of Lemma 12 is not true even for very nice spaces; as shown in [10, example 4C]. More precisely, put $E=F=L^{2}[0,1]$. Then, the Fremlin projective tensor product $E\widehat{\otimes}F$ is not order complete so that the projective norm on $E\widehat{\otimes}F$ is not order continuous, although both $E$ and $F$ are order continuous (even reflexive) Banach lattices.
Proposition 13.
Suppose $E$ and $F$ are Banach lattices. Assume that $(x_{\alpha})\subseteq E$ is $un$-null and $(y_{\alpha})\subseteq F$ is eventually norm bounded. Then, $x_{\alpha}\otimes y_{\alpha}\xrightarrow{un}0$ in $E\widehat{\otimes}F$.
Proof.
By the assumption, $\|y_{\alpha}\|\leq 1$ for sufficiently large $\alpha$. Suppose $w\in(E\overline{\otimes}F)_{+}$. By [10, 1A (d)]), there exist $x_{0}\in E_{+}$ and $y_{0}\in F_{+}$ with $w\leq x_{0}\otimes y_{0}$. Now, we use the following inequality.
$|x_{\alpha}\otimes y_{\alpha}|\wedge w\leq|x_{\alpha}|\otimes|y_{\alpha}|\wedge x_{0}\otimes y_{0}$
$\leq(|x_{\alpha}|\wedge x_{0})\otimes(|y_{\alpha}|\vee y_{0}).$
Therefore,
$\||x_{\alpha}\otimes y_{\alpha}|\wedge w\|\leq\|\leq\||x_{\alpha}|\wedge x_{0}\|\||y_{\alpha}|\vee y_{0}\|\rightarrow 0.$
We conclude that $x_{\alpha}\otimes y_{\alpha}\xrightarrow{un}0$ in $E\overline{\otimes}F$ and so that in $E\widehat{\otimes}F$ by [14, Theorem 4.3]; see also [16, Lemma 3.5]. ∎
Remark 14.
Note that norm boundedness is essential in Proposition 13 and can not be omitted. Put $E=\ell_{\infty}$. Consider two sequences $(u_{n})$ and $(v_{n})$ in $E$ defined via $u_{n}=(0,\ldots,0,n,0,\ldots)$ and $v_{n}=(0,\dots,0,\frac{1}{n},0,\ldots)$. Indeed, $v_{n}\xrightarrow{un}0$ and $(u_{n})$ is unbounded. Consider the Fremlin tensor product $E\widehat{\otimes}E$. The sequence $(u_{n}\otimes v_{n})_{n\in\mathbb{N}}$ is the sequence $(e_{n})$ consisting of the standard basis of $E$; that is not $un$-null by [14, Theorem 3.2] (note that $E\widehat{\otimes}E$ is a Banach lattice with order unit).
Theorem 15.
Suppose $E$ and $F$ are Banach lattices such that $E\widehat{\otimes}F$ is order continuous. If $(x_{\alpha})\subseteq E$ is $un$-convergent to $x\in E$ and $(y_{\alpha})\subseteq F$ is $un$-convergent to $y\in F$, then $x_{\alpha}\otimes y_{\alpha}\xrightarrow{un}x\otimes y$ in $E\widehat{\otimes}F$.
Proof.
First note that by Lemma 12, both $E$ ad $F$ are order continuous. Moreover, observe that by [8, Corollary 3.5], there exists an increasing sequence $(\alpha_{k}^{\prime})$ of indices such that $x_{\alpha_{k}^{\prime}}-x\xrightarrow{un}0$. By [8, Theorem 3.2], there exists a subsequence $(\alpha_{k})$ of $(\alpha_{k}^{\prime})$ and a disjoint sequence $(d_{k})$ in $E$ such that $\|x_{\alpha_{k}}-x-d_{k}\|\rightarrow 0$. By Proposition 13, $(x_{\alpha_{k}}-x-d_{k})\otimes(y_{\alpha}-y)\xrightarrow{un}0$, $(x_{\alpha_{k}}-x)\otimes y\xrightarrow{un}0$ and $x\otimes(y_{\alpha}-y)\xrightarrow{un}0$. Furthermore, observe that the net $d_{k}\otimes(y_{\alpha}-y)$ is disjoint in $E\widehat{\otimes}F$ so that $un$-null by [14, proposition 3.5]. Therefore, we have
$x_{\alpha_{k}}\otimes y_{\alpha}-x\otimes y=(x_{\alpha_{k}}-x-d_{k})\otimes(y_{\alpha}-y)+d_{k}\otimes(y_{\alpha}-y)+$
$x\otimes(y_{\alpha}-y)+(x_{\alpha_{k}}-x)\otimes y\xrightarrow{un}0.$
This procedure happens for every subnet of $(x_{\alpha})$ and $(y_{\alpha})$; since $un$-convergence is topological, we conclude that $x_{\alpha}\otimes y_{\alpha}\xrightarrow{un}x\otimes y$. ∎
Question 3.
Does the result of Theorem 15 hold without order continuity of the Fremlin projective tensor product $E\widehat{\otimes}F$.
Now, we investigate the relation between unbounded absolute weak convergence and the Fremlin projective tensor product.
Proposition 16.
Suppose $E$ and $F$ are Banach lattices. Assume that $(x_{\alpha})\subseteq E$ is $uaw$-null and $(y_{\alpha})\subseteq F$ is eventually order bounded. Then, $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uaw}0$ in $E\widehat{\otimes}F$.
Proof.
By the assumption, there exists $y_{1}\in F_{+}$ such that $y_{\alpha}\leq y_{1}$ for sufficiently large $\alpha$. Suppose $w\in(E\overline{\otimes}F)_{+}$. By [10, 1A (d)]), there exist $x_{0}\in E_{+}$ and $y_{0}\in F_{+}$ with $w\leq x_{0}\otimes y_{0}$. Fix $f\in(E\widehat{\otimes}F)^{{}^{\prime}}_{+}=B^{r}(E\times F)_{+}$, where $B^{r}(E\times F)$ is the Banach lattice of all bounded bilinear forms defined on $E\times F$; see [10, 1I] for more details. Now, we use the following inequality.
$|f(x_{\alpha}\otimes y_{\alpha}\wedge w)|\leq f(|x_{\alpha}|\otimes|y_{\alpha}|\wedge x_{0}\otimes y_{0})$
$\leq f(|x_{\alpha}|\wedge x_{0},|y_{\alpha}|\vee y_{0})\leq f(|x_{\alpha}|\wedge x_{0},y_{0}\vee y_{1}).$
Note that the restriction $f_{y_{0}\vee y_{1}}:E\to\mathbb{R}$ is a positive linear functional. Therefore, by the assumption,
$f(|x_{\alpha}\otimes y_{\alpha}|\wedge w)\leq f(|x_{\alpha}\wedge x_{0},y_{0}\vee y_{1})\rightarrow 0.$
We conclude that $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uaw}0$ in $E\overline{\otimes}F$ and so that in $E\widehat{\otimes}F$ by [16, Lemma 3.4]. ∎
Theorem 17.
Suppose $E$ and $F$ are Banach lattices such that of one them is order continuous. If $(x_{\alpha})\subseteq E$ is $uaw$-convergent to $x\in E$ and $(y_{\alpha})\subseteq F$ is $uaw$-convergent to $y\in F$, then $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uaw}x\otimes y$ in $E\widehat{\otimes}F$.
Proof.
Assume that $E$ is order continuous. By [19, Theorem 4], $x_{\alpha}\xrightarrow{un}x$. By [8, Theorem 3.2], there exists a increasing sequence $(\alpha_{k})$ of indices and a disjoint sequence $(d_{k})$ in $E$ such that $\|x_{\alpha_{k}}-x-d_{k}\|\rightarrow 0$. By passing to a further subsequence, we may assume that $(x_{\alpha_{k}}-x-d_{k})$ is order bounded. So, by Proposition 16, $(x_{\alpha_{k}}-x-d_{k})\otimes(y_{\alpha}-y)\xrightarrow{uaw}0$, $(x_{\alpha_{k}}-x)\otimes y\xrightarrow{uaw}0$ and $x\otimes(y_{\alpha}-y)\xrightarrow{uaw}0$. Note that the net $(d_{k}\otimes(y_{\alpha}-y))$ is disjoint in $E\widehat{\otimes}F$ so that $uaw$-null by [19, Lemma 2]. Thus, we conclude that
$x_{\alpha_{k}}\otimes y_{\alpha}-x\otimes y=(x_{\alpha_{k}}-x-d_{k})\otimes(y_{\alpha}-y)+d_{k}\otimes(y_{\alpha}-y)+$
$x\otimes(y_{\alpha}-y)+(x_{\alpha_{k}}-x)\otimes y\xrightarrow{uaw}0.$
This argument happens for every subnet; therefore, $x_{\alpha}\otimes y_{\alpha}\xrightarrow{uaw}x\otimes y$. ∎
Question 4.
Is order continuity of one of the components necessary as an assumption for Theorem 17?
@@ 3.2. Fremlin projective tensor product and unbounded convergences
Acknowledgments. Most of the work on this paper was done during a visit of the author to University of Alberta in 2023. This note would not have existed without invaluable help and discussions of V. G. Troitsky, my friend and my colleague. The author would like to thank him so much for great hospitality during his visit, as well. Thanks is also due to Eugene Bilokopytov for many useful discussions.

### v4

@@ 1. Introduction
By the known Maeda-Ogasawara theorem ([1, Theorem 7.29]) and using Lemma 1, There are two compact Hausdorff extremally disconnected topological spaces $\Sigma$ and $\Omega$ such that $E$ is an order dense vector sublattice of $S(\Sigma)$ and $F$ is an order dense vector sublattice of $S(\Omega)$. Thus, by [1, Theorem 1.23], $E$ and $F$ are regular sublattices in $S(\Sigma)$ and $S(\Omega)$, respectively. So, by [9, Theorem 3.2], $f_{\alpha}\xrightarrow{uo}f$ in $S(\Sigma)$ and $g_{\beta}\xrightarrow{uo}g$ in $S(\Omega)$. On the other hand, by [5, Proposition 3.1], $E\overline{\otimes}F$ can be considered as a vector sublattice of $S(\Sigma\times\Omega)$ generated by the mappings $(h\otimes k)(t,s)=h(t)k(s)$ for each $h\in S(\Sigma)$ and for each $k\in S(\Omega)$. Compatible with [4, Remark 4.1] and Lemma 1, it is enough to show that $f_{\alpha}\otimes g_{\beta}\xrightarrow{uo}f\otimes g$ in $S(\Sigma\times\Omega)$.
Now, we use Lemma 4. Suppose $W\subseteq\Sigma\times\Omega$ is a non-empty open set and $\varepsilon\in(0,1)$ is arbitrary. There are some non-empty open sets $U\subseteq\Sigma$ and $V\subseteq\Omega$ such that $U\times V\subseteq W$. We can find non-empty open sets $U_{1}\subseteq U$ and $V_{1}\subseteq V$ such that $|f|$ is uniformly bounded (say, by $K_{1}$) on $U_{1}$ and $|g|$ is uniformly bounded on $V_{1}$ (namely, by $K_{2}$). Choose open subsets $U_{2}\subseteq U_{1}$ and $V_{2}\subseteq V_{1}$ with $|(f_{\alpha}-f)(t)|<\frac{\varepsilon}{K_{1}}$ for each $t\in U_{2}\cap Dom(f_{\alpha})\cap Dom(f)$ and $|(g_{\beta}-g)(s)|<\frac{\varepsilon}{K_{2}}$ for each $s\in V_{2}\cap Dom(g_{\beta})\cap Dom(g)$, provided that $\alpha$ and $\beta$ are sufficiently large.
Then for each $(t,s)\in(U_{2}\times V_{2})\cap((Dom(f_{\alpha})\cap Dom(f))\times(Dom(g_{\beta})\cap Dom(g)))=(U_{2}\cap Dom(f_{\alpha})\cap Dom(f))\times(V_{2}\cap Dom(g_{\beta})\cap Dom(g)))$ and for sufficiently large $\alpha$ and $\beta$, we have
$|(f_{\alpha}\otimes g_{\beta}-f\otimes g)(t,s)|\leq|(f_{\alpha}\otimes(g_{\beta}-g))(t,s)|+|((f_{\alpha}-f)\otimes g)(t,s)|$
$=|f_{\alpha}(t)||(g_{\beta}-g)(s)|+|(f_{\alpha}-f)(t)||g(s)|<(\frac{K_{1}}{K_{2}}+\frac{K_{2}}{K_{1}}+\frac{\varepsilon}{K_{1}K_{2}})\varepsilon.$
@@
Now, an easy application of Lemma 5 and Lemma 6 convinces us that $f_{\alpha}\otimes g_{\beta}\xrightarrow{uo}f\otimes g$ in $E\overline{\otimes}F$.
@@
Suppose $E$ and $F$ are vector lattices. Assume that $(x_{\alpha})\subseteq E$ is $uo$-null and $(y_{\beta})\subseteq F$ is eventually order bounded. Then, $x_{\alpha}\otimes y_{\beta}\xrightarrow{uo}0$ in $E\overline{\otimes}F$.
@@
There exists $y\in F_{+}$ such that $|y_{\beta}|\leq y$ for sufficiently large $\beta$. Suppose $w\in(E\overline{\otimes}F)_{+}$. By [8, 1A (d)]), there exist $x_{0}\in E_{+}$ and $y_{0}\in F_{+}$ with $w\leq x_{0}\otimes y_{0}$. Now, we use the following inequality.
$|x_{\alpha}\otimes y_{\beta}|\wedge w\leq(|x_{\alpha}|\otimes|y_{\beta}|)\wedge(x_{0}\otimes y_{0})$
$\leq(|x_{\alpha}|\wedge x_{0})\otimes(|y_{\beta}|\vee y_{0})\leq(|x_{\alpha}|\wedge x_{0})\otimes(y\vee y_{0}).$
Now, by [10, Corollary 3.4], we conclude that $x_{\alpha}\otimes y_{\beta}\xrightarrow{uo}0$, as claimed. ∎
@@
Note that order bounded assumption is essential in Proposition 8 and can not be removed. It is know that in $L_{p}(\mu)$-spaces, for sequences, $uo$-convergence and almost everywhere convergence agree. Put $E=L^{2}(\mathbb{R})$, $f_{n}(x)=\frac{x}{n}$ and $g_{n}(x)=n$. It is easy to see that $(f_{n})$ is $uo$-null and $(g_{n})$ is not order bounded. Moreover, $(f_{n}\otimes g_{n})_{n}$ is the identity mapping which is not pointwise convergent to zero, certainly.
It can be easily verified that if a net $(x_{\alpha})\subseteq E$ and a net $(y_{\beta})\subseteq F$ are eventually order bounded, then, the net $(x_{\alpha}\otimes y_{\beta})$ is also eventually order bounded in $E\overline{\otimes}F$. So, we have the following fact that is an extension of [10, Corollary 3.4], as well.
@@
Acknowledgments. Most of the work on this paper was done during a visit of the author to University of Alberta in 2023. The author would like to thank Professor Vladimir Troitsky so much for great hospitality during his visit, as well. Thanks is also due to Eugene Bilokopytov for many useful discussions.